A student in the fitness center lifts a weight with m = 68.5 kg from the floor to a height of 1.90 meters in a time of 13.5 seconds. Find the average power exerted by the student during this time. Give your answer in watts (W) and with 3 significant figures.
Answer #1

Solution:

Given:

m = 68.5 kg

h = 1.90 m

t = 13.5 sec

According to the formula for average power is :

P = W / t = (m g h) / t = { (68.5 kg)( 9.8 m/s2 )( 1.90 m) } / (13.5 s) = 94.5 W

Therefore: the average power exerted by the student during this time is : P = 94.5 W